I 4. An alternating voltage source (15 V rms) drives the circuit shown here at f=1.00 kHz, where L = 175 mH and R= 1.80 k22,Need help solving part d.
Answer #1

Halas X, - 270+2 = (270)(109H2) (0-175H) - to 11002 R = 18002 2. R++*+ = 2102 Ans (0) ® Speak - B Vime : 0.01014 10.164 TF O